Avoiding a collision.
A car is traveling on a road that
is perpendicular to a railroad track. When the car is
30 meters from the crossing, the car’s new collision
detector warns the driver that there is a train 50 meters
from the car and heading toward the same crossing. How
far is the train from the crossing?

Answer by mananth(16946)

the car , the crossing and the train at any given point are at right angles till they meet.
car - train distance = 50 meters
car-crossing distance = 30 meters
so we have to find the other leg of the right triangle.
50^2-30^2= d^2
2500-900=d^2
1600=d^2
d= 40 meters distance of the train from the crossing.
...
m.ananth@hotmail.ca